# Adaptively Bypassing Vision Transformer Blocks for Efficient Visual Tracking

### v1

@@ 3.3 Vision Transformer Pruning (VTP)
While the architecture of the proposed BDM is simple, the computational burden it introduces cannot be overlooked. To combat this, we adapt the ViT pruning method proposed in Zhu et al. (2021) to mitigate the additional time cost and enhance efficiency. Our enhancements extend the applicability of the original method, formulated for scaled dot-product attention, to ViTs based on multi-headed self-attention (MSA). Additionally, we employ a local ranking of importance scores within each Vision Transformer (ViT) block, departing from the original global approach. These adaptations not only make our proposed method applicable to more commonly used ViTs but also overcome the weakness inherent in the global ranking of importance scores. In the original approach, the average scores across all ViT blocks are ranked to determine the pruning threshold. However, this method may lead to under-pruning in ViT layers with relatively higher overall importance scores, while layers with relatively smaller values may undergo over-pruning. The proposed ViT pruning aims to evaluate the impact of dimensions in each layer and selectively prunes those dimensions with low importance scores within each block locally. Let $\mathcal{X}\in\mathbb{R}^{d\times n}$ denote an input sequence of $n$ tokens of dimensions $d$, the proposed pruning-enabled ViT layer that map $\mathcal{X}$ to $\mathcal{Z}$ is formualted as follows,
$\begin{split}&\mathcal{Y}=\mathbf{\check{MSA}}(\mathbf{LN}(\mathcal{X}))+\mathcal{X},\\ &\mathcal{Z}=\mathbf{\check{MLP}}(\mathbf{LN}(\mathcal{Y}))+\mathcal{Y},\end{split}$ (3)
@@
where $\mathbf{LN}$ denotes layer normalization, $\mathbf{\check{MSA}}$ and $\mathbf{\check{MLP}}$ are generalized MSA and multi-layer perceptron (MLP). Suppose $\mathbf{\check{MSA}}$ is parameterized by $\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V},\textbf{W}^{O}\in\mathbb{R}^{d\times d}$, $N_{h}\in\mathbb{N}$, i.e., number of heads, $\mathbf{\check{MLP}}$ is parameterized by $\textbf{W}^{\mathcal{K}}\in\mathbb{R}^{d\times d},\mathcal{K}=1,...,N_{l},N_{l}\in\mathbb{N}$, i.e., number of linear layers, then they are formally defined as follows [^1] Other parameters are omitted for simplification.,
@@
$\begin{split}&\mathbf{\check{MSA}}(\mathcal{X};\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V},\textbf{W}^{O},N_{h})\\ &=Concat(\{\mathbf{Attention}(\widetilde{\mathcal{X}}\widetilde{\textbf{W}}^{Q}_{i},\widetilde{\mathcal{X}}\widetilde{\textbf{W}}^{K}_{i},\widetilde{\mathcal{X}}\widetilde{\textbf{W}}^{V}_{i})\}^{N_{h}}_{i=1})\widetilde{\textbf{W}}^{O},\\ &\mathbf{\check{MLP}}(\mathcal{Y};\textbf{W}^{1},...,\textbf{W}^{{N_{l}}})=\phi_{{N_{l}}}(...(\phi_{1}(\widetilde{Y}\widetilde{\textbf{W}}^{1})...)\widetilde{\textbf{W}}^{{N_{l}}}),\end{split}$ (4)
@@
where $Concat$ performs the concatenation operation, and
@@
$\begin{split}&\textbf{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\textbf{Softmax}(\textbf{K}^{T}\textbf{Q}/\sqrt{d})\textbf{V},\\ &\textbf{W}^{\mathcal{J}}=[\textbf{W}^{\mathcal{J}}_{1},...,\textbf{W}^{\mathcal{J}}_{N_{h}}]\quad s.t.\quad\mathcal{J}\in\{Q,K,V\},\\ &\widetilde{\textbf{W}}^{\mathcal{I}}=\textbf{D}_{1}{\textbf{W}}^{\mathcal{I}}\quad s.t.\quad\mathcal{I}\in\{Q,K,V,O\},\quad\widetilde{\mathcal{X}}=\mathcal{X}\textbf{D}_{1},\\ &\widetilde{\textbf{W}}^{\mathcal{I}}=\textbf{D}_{2}{\textbf{W}}^{\mathcal{K}}\quad s.t.\quad\mathcal{K}\in\{1,...,N_{l}\},\quad\widetilde{\mathcal{Y}}=\mathcal{Y}\textbf{D}_{2}.\end{split}$ (5)
@@
$\textbf{Softmax}(\cdot)$ denotes the row-wise softmax normalization function, $\phi_{i}$ denotes the activation function at layer $i$ of $\mathbf{\check{MLP}}$. $\textbf{D}_{1},\textbf{D}_{2}\in\mathfrak{D}:=\{diag(\mathbf{v})|\mathbf{v}\in\{0,1\}^{d},N_{h}\mid||\mathbf{v}||_{1}\}$ are two diagonal matrices used to select or mask out dimensions, called dimension reduction (DR) matrices. If $\textbf{D}_{1}$ and $\textbf{D}_{2}$ are identity matrix, then $\mathbf{\check{MSA}}$ and $\mathbf{\check{MLP}}$ reduce to the traditional formulations of Multi-Head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) as presented in prior works  Dosovitskiy et al. (2021); Vaswani et al. (2017). It is important to note that, unlike in Zhu et al. (2021), in our setting, the $L_{1}$ norm $||\mathbf{v}||_{1}$ must be divisible by $N_{h}$ due to the consideration of the MSA setting. It’s worth noting that $\mathbf{\check{MSA}}$ and $\mathbf{\check{MLP}}$ are formally identical to their traditional counterparts.
@@
The objective of our ViT pruning is the same as Zhu et al. (2021), namely, $\textbf{D}_{1}$ and $\textbf{D}_{2}$ are optimized as trainable parameters with back-propagation algorithm for each ViT layer. Considering difficulties due to their discrete values, they are relaxed to real values for end-to-end optimization. Let $\hat{\textbf{D}}_{1}^{k},\hat{\textbf{D}}_{2}^{k}\in\hat{\mathfrak{D}}:=\{diag(\mathbf{v})|\mathbf{v}\in\mathbb{R}^{d}\}$ denote the relaxed DR matrices at layer $k$ of the ViT to be pruned. To enforce sparsity of all DR matrices and thus achieve a high model pruning ratio, a $L_{1}$ regularization term is added to the training loss, which is
@@
where $\alpha_{k}$ is a sparsity hyper-parameter to control the sparsity at layer $k$ and is set to $10^{-4}$ by default. After the training for ViT pruning, $\{\hat{\textbf{D}}_{1}^{k},\hat{\textbf{D}}_{2}^{k}\}_{k=1}^{N}$ are binarized to obtain the DR matrices according to a predefined pruning ratio $\mu$, namely, by setting the first $d^{*}=\lfloor d\times\sigma/N_{h}\rfloor\times N_{h}$ values ranked in descending order to ones and the rest to zeros. The pipeline for pruning is summarized as follows: 1) traning with the sparsity regularization; 2) pruning dimensions of the parameter matrices with DR matrices; 3) fine-tuning.
@@ 3.4 Prediction Head and Training Objective
As in Cui et al. (2022b); Ye et al. (2022b), we employ a prediction head that consists of multiple Conv-BN-ReLU layers, to directly estimate the bounding box of the target. The output tokens associated with the search image are initially transformed into a 2D spatial feature map before being fed into the prediction head. The prediction head produces a target classification score $\mathbf{p}\in[0,1]^{H_{x}/P\times W_{x}/P}$, a local offset $\mathbf{o}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$, and a normalized bounding box size $\mathbf{s}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$, $H_{x}$ and $W_{x}$ respectively represent the height and width of the search image, while $P$ represents the patch size. The initial estimate of the target position is determined by identifying the location with the highest classification score, i.e., $(x_{c},y_{c})=\textup{argmax}_{(x,y)}\mathbf{p}(x,y)$, based on which the final target bounding box is estimated by
@@
For the tracking task, we adopt the weighted focal loss $\mathcal{L}_{cls}$ Law and Deng (2018) for classification, a combination of $L_{1}$ loss and GIoU loss $\mathcal{L}_{iou}$ Rezatofighi et al. (2019) for bounding box regression. Finally, the overall loss function is:
@@
where the constants $\lambda_{iou}=2$ and $\lambda_{L_{1}}=5$ are the same as in Cui et al. (2022b); Ye et al. (2022b), $\gamma$ is set to 5. Our framework is trained end-to-end with the overall loss $\mathcal{L}_{overall}$ after the pretrained weights of the ViT for image classification is loaded.
Table 1: Compared with the state-of-the-art methods on the GOT-10k Huang et al. (2021), LaSOT Fan et al. (2018), TrackingNet Mueller et al. (2018) and UAV123 Mueller et al. (2016) benchmarks. The top three real-time results are displayed in Red, blue and green fonts, while the best non-real-time results are in bold font.
Method Source GOT-10K LaSOT TrackingNet UAV123 Speed(FPS)
$AO$ $SR_{0.5}$ $SR_{0.75}$ $AUC$ $P_{Norm}$ $P$ $AUC$ $P_{Norm}$ $P$ $AUC$ $P$ GPU CPU AGX
ABTrack-DeiT 64.0 74.2 56.1 61.1 70.1 62.8 78.1 83.2 74.3 64.4 81.8 223 64 71
ABTrack-ViT 61.9 72.2 53.0 59.9 68.8 60.8 77.2 82.5 73.1 64.3 81.9 241 69 73
MixFormerV2 Cui et al. (2023) NIPS23 61.9 71.7 51.3 60.6 69.9 60.4 75.8 81.1 70.4 63.7 81.3 167 45 49
HiT Kang et al. (2023) ICCV23 62.6 71.2 54.4 60.5 68.3 61.5 77.7 81.9 73.1 63.3 80.5 173 47 59
E.T.Track Blatter et al. (2021) WACV23 62.5 71.4 55.9 59.1 68.2 60.5 72.5 77.8 69.5 62.3 81.5 56 32 22
FEAR Borsuk et al. (2022) ECCV22 61.9 72.2 52.5 53.5 59.7 54.5 70.2 80.8 71.5 58.9 79.2 119 34 38
TCTrack Cao et al. (2022) CVPR22 66.2 75.6 61.0 60.5 69.3 62.4 74.8 79.6 73.3 60.5 80.0 140 45 41
LightTrack Yan et al. (2021b) CVPR21 61.1 71.0 54.3 53.8 60.5 53.7 72.5 77.8 69.5 59.9 77.6 134 39 36
Non-real-time ARTrack  Wei et al. (2023a) CVPR23 73.5 82.2 70.9 70.4 79.5 76.6 84.2 88.7 83.5 67.9 85.9 26 9 8
GRM Gao et al. (2023) CVPR23 73.4 82.9 70.4 69.9 79.3 75.8 84.0 88.7 83.3 67.7 85.4 53 18 16
CTTrack Song et al. (2023) AAAI23 71.3 80.7 70.3 67.8 77.8 74.0 82.5 87.1 80.3 68.8 89.5 40 13 12
OSTrack Ye et al. (2022a) ECCV22 70.5 80.9 66.7 69.1 78.7 75.2 81.9 87.2 81.0 67.5 84.7 68 18 19
CSWinTT Song et al. (2022) CVPR22 69.4 78.9 65.4 66.2 75.2 70.9 81.9 86.7 79.5 70.5 90.3 10 3 3
ToMP Mayer et al. (2022) CVPR22 68.1 78.3 61.3 68.5 79.2 73.5 81.5 86.4 78.9 68.5 87.3 24 8 7
TrSiam Wang et al. (2021a) CVPR21 67.3 78.7 58.6 62.4 70.5 60.6 78.1 82.9 72.7 67.4 85.3 55 7 10
@@ 4 Experiments
In this section, we extensively evaluate our method on five tracking benchmarks, including comparison with SOTA methods, comprehensive ablation and generality study.
@@ 4.1 Implementation Details
Model. Our method is a general approach based on Vision Transformer (ViT) models.
In this paper, we adopt three efficient ViT models, i.e., A-ViT-T Yin et al. (2022), DeiT-Tiny Touvron et al. (2021), and ViT-Tiny Dosovitskiy et al. (2020), as the backbones to develop three efficient trackers, named ABTrack-AViT, ABTrack-DeiT, and ABTrack-ViT, respectively. All our models are implemented with Python 3.8 and PyTorch 1.9.0.
Training. The training data comprise four datasets, i.e., the combination of the training sets of GOT-10k Huang et al. (2021), LaSOT Fan et al. (2018), COCO Lin et al. (2014), and TrackingNet Mueller et al. (2018). The batch size is set to 32. We set the bypass probability threshold $\rho$, block sparsity parameter $\tau_{0}$, $\varrho$, and pruning ratio $\mu$ to 0.5, 0.4, 0.1, and 0.3, respectively. We use the AdamW optimizer Loshchilov and Hutter (2017) with a weight decay of $10^{-4}$ for finally training the model, the initial learning rate of the backbone is set to $4\times 10^{-5}$, and the models are trained for 300 epochs with 60k image pairs per epoch. The learning rate is reduced by a factor of 10 after 240 epochs. We use four NVIDIA P100 GPUs for training.
Inference. During the inference phase, we apply a Hann window penalty as a position prior to the tracked object, following the conventional practice Zhang and Peng (2020). This penalty serves to enforce spatial constraints during the tracking process, thereby enhancing tracking accuracy.
@@ 4.2 Datasets
GOT-10k Huang et al. (2021) is a large-scale benchmark for generic object tracking in the wild. It contains over 10,000 video sequences with diverse challenges such as occlusion, motion blur, and scale variations. TrackingNet Mueller et al. (2018) is a high-quality dataset and benchmark for object tracking. It consists of 30,000 video sequences with precise pixel-level annotations, covering a wide range of object categories and tracking scenarios. LaSOT Fan et al. (2018) is another comprehensive benchmark for single object tracking. It contains 1,400 video sequences with a total of 3.5 million frames, covering various challenges like scale variations, occlusions, and camera motion. UAV123 Mueller et al. (2016) is a dataset specifically designed for UAV tracking. It consists of 123 video sequences captured by drones, featuring challenging factors such as fast motion, low resolution, and viewpoint changes. VOT-ST2021 Kristan et al. (2021) is a benchmark designed to evaluate the performance of short-term visual object tracking algorithms. It serves as a platform for researchers and developers to assess the capabilities of various tracking methods in real-world scenarios

### v2

@@ 3.3 Vision Transformer Pruning (VTP)
While the architecture of our proposed Bypass Decision Module (BDM) is straightforward, it introduces a computational burden that cannot be ignored. To tackle this issue, we refine the ViT pruning method introduced in [20] to minimize additional time costs and enhance efficiency. Our enhancements broaden the scope of the original method, initially designed for scaled dot-product attention mechanisms, to encompass ViTs employing multi-headed self-attention (MSA). Additionally, we employ a local ranking of importance scores within each ViT block, departing from the original global approach, as demonstrated in Figure 3. In the original approach, as shown in Figure 3 (a), pruning thresholds were determined based on the average scores across all ViT blocks. However, this method could lead to under-pruning in layers with higher overall importance scores and over-pruning in layers with lower scores. Our proposed ViT pruning method, as shown in Figure 3 (b), evaluates the impact of dimensions within each layer locally, selectively pruning dimensions with low importance scores in each block. These improvements not only extend the applicability of our proposed method to widely used ViTs but also address the limitations associated with the global ranking of importance scores, which are detailed as follows.
(a) global ranking pruning
(b) local ranking pruning
Figure 3: Comparison between global ranking pruning and local ranking pruning, where the grey blocks represent the dimensions pruned in each transformer block.
Let $\mathcal{X}\in\mathbb{R}^{d\times n}$ denote an input sequence of $n$ tokens of dimensions $d$, the proposed pruning-enabled ViT layer that map $\mathcal{X}$ to $\mathcal{Z}$ is defined as follows,
$\small\begin{split}&\mathcal{Y}=\mathbf{\check{MSA}}(\mathbf{LN}(\mathcal{X}))+\mathcal{X},\\ &\mathcal{Z}=\mathbf{\check{MLP}}(\mathbf{LN}(\mathcal{Y}))+\mathcal{Y},\end{split}$ (3)
@@
where $\mathbf{LN}$ denotes layer normalization, $\mathbf{\check{MSA}}$ and $\mathbf{\check{MLP}}$ are generalized MSA and multi-layer perceptron (MLP). Suppose $\mathbf{\check{MSA}}$ is parameterized by $\textbf{W}_{i}^{Q},\textbf{W}_{i}^{K},\textbf{W}_{i}^{V}\in\mathbb{R}^{d\times(d/N_{h})}$, $\textbf{W}^{O}\in\mathbb{R}^{d\times d}$, for $i=1,...,N_{h}$, and $N_{h}\in\mathbb{N}$ (i.e., the number of heads that divides $d$). Additionally, let $\mathbf{\check{MLP}}$ be parameterized by $\textbf{W}_{k}^{L}\in\mathbb{R}^{d\times d}$ for $k=1,...,N_{l}$, where $N_{l}\in\mathbb{N}$ represents the number of linear layers. $\mathbf{\check{MSA}}$ and $\mathbf{\check{MLP}}$ are formally defined as follows,
@@
$\small\begin{split}&\mathbf{\check{MSA}}(\mathcal{X};\{\textbf{W}^{Q},\textbf{W}_{i}^{K},\textbf{W}_{i}^{V}\}^{N_{h}}_{i=1},\textbf{W}^{O})\\ &=\text{Concat}(\{\mathbf{Attention}(\widetilde{\mathcal{X}}\widetilde{\textbf{W}}^{Q}_{i},\widetilde{\mathcal{X}}\widetilde{\textbf{W}}^{K}_{i},\widetilde{\mathcal{X}}\widetilde{\textbf{W}}^{V}_{i})\}^{N_{h}}_{i=1})\widetilde{\textbf{W}}^{O},\\ &\mathbf{\check{MLP}}(\mathcal{Y};\textbf{W}^{L}_{1},...,\textbf{W}^{L}_{{N_{l}}})=\phi_{{N_{l}}}(...(\phi_{1}(\widetilde{Y}\widetilde{\textbf{W}}^{L}_{1})...)\widetilde{\textbf{W}}^{L}_{{N_{l}}}),\end{split}$ (4)
@@
where Concat performs the concatenation operation, $\widetilde{\mathcal{X}}=\mathcal{X}\textbf{D}_{1}$, $\widetilde{\textbf{W}}^{O}=\textbf{D}_{1}{\textbf{W}}^{O}$, $\widetilde{\mathcal{Y}}=\mathcal{Y}\textbf{D}_{2}$, and
@@
$\small\begin{split}&\textbf{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\textbf{Softmax}(\textbf{K}^{T}\textbf{Q}/\sqrt{d})\textbf{V},\\ &\widetilde{\textbf{W}}_{i}^{J}=\textbf{D}_{1}{\textbf{W}}_{i}^{J}\quad\forall\quad J\in\{Q,K,V\}\quad\&\quad i\in\{1,...,N_{h}\}\\ &\widetilde{\textbf{W}}_{k}^{L}=\textbf{D}_{2}{\textbf{W}}_{k}^{L}\quad\forall\quad k\in\{1,...,N_{l}\},\quad\end{split}$ (5)
@@
in which $\textbf{Softmax}(\cdot)$ denotes the row-wise softmax normalization function, $\phi_{i}$ denotes the activation function at layer $i$ of $\mathbf{\check{MLP}}$, and $\textbf{D}_{1},\textbf{D}_{2}\in\mathfrak{D}:=\{diag(\mathbf{v})|\mathbf{v}\in\{0,1\}^{d},N_{h}\mid||\mathbf{v}||_{1}\}$ are two diagonal matrices used to select or mask out dimensions, called dimension reduction (DR) matrices. $||\cdot||_{1}$ indicates the $l^{1}$ norm. If $\textbf{D}_{1}$ and $\textbf{D}_{2}$ are identity matrix, then $\mathbf{\check{MSA}}$ and $\mathbf{\check{MLP}}$ reduce to the traditional formulations of Multi-Head Self-Attention (MSA) and Multi-Layer Perceptron (MLP) as presented in prior works  [15, 5].
@@
The objective of our proposed ViT pruning method is to optimize $\textbf{D}_{1}$ and $\textbf{D}_{2}$ as trainable parameters for each ViT layer. By treating these matrices as trainable, we can dynamically adjust the pruning process during training, allowing for a more fine-tuned and effective reduction of computational complexity. This approach leverages the inherent flexibility of deep learning optimization techniques, enabling the model to learn which dimensions are most critical for preserving performance while minimizing unnecessary computations. Considering optimization difficulties due to their discrete values, $\textbf{D}_{1}$ and $\textbf{D}_{2}$ are relaxed to real values for end-to-end optimization. Let $\hat{\textbf{D}}_{1}^{k},\hat{\textbf{D}}_{2}^{k}\in\hat{\mathfrak{D}}:=\{diag(\mathbf{v})|\mathbf{v}\in\mathbb{R}^{d}\}$ denote the relaxed DR matrices at layer $k$ of the ViT to be pruned. To enforce sparsity of all DR matrices and thereby achieve a higher model pruning ratio, a $L_{1}$ regularization term is added to the training loss, which is
@@
where $\alpha_{k}$ is a sparsity hyper-parameter to control the sparsity at layer $k$. After the ViT pruning training, the relaxed DR matrices $\{\hat{\textbf{D}}_{1}^{k},\hat{\textbf{D}}_{2}^{k}\}_{k=1}^{N}$ are binarized according to a predefined pruning ratio $\mu$. This process involves setting the first $d^{*}=\lfloor d\times\mu/N_{h}\rfloor\times N_{h}$ values, ranked in descending order, to ones, and the remaining values to zeros. This binarization step ensures that the pruning is consistent with the desired level of sparsity, enabling efficient model performance while retaining the most critical features. The pipeline for pruning is summarized as follows: 1) training with the sparsity regularization; 2) pruning dimensions of the parameter matrices with DR matrices; 3) fine-tuning.
@@ 3.4 Prediction Head and Training Objective
As in [9, 10], we employ a prediction head that consists of multiple Conv-BN-ReLU layers, to directly estimate the bounding box of the target. The output tokens associated with the search image are initially transformed into a 2D spatial feature map before being fed into the prediction head. The prediction head produces a target classification score $\mathbf{p}\in[0,1]^{H_{x}/P\times W_{x}/P}$, a local offset $\mathbf{o}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$, and a normalized bounding box size $\mathbf{s}\in[0,1]^{2\times H_{x}/P\times W_{x}/P}$, where $H_{x}$ and $W_{x}$ respectively represent the height and width of the search image, while $P$ represents the patch size. The initial estimate of the target position is determined by identifying the location with the highest classification score, i.e., $(x_{c},y_{c})=\textup{argmax}_{(x,y)}\mathbf{p}(x,y)$, based upon which the final target bounding box is estimated by
@@
For the tracking task, we adopt the weighted focal loss $\mathcal{L}_{cls}$ [43] for classification, a combination of the $L_{1}$ loss and the GIoU loss $\mathcal{L}_{iou}$ [42] for bounding box regression. Finally, the overall loss function is:
@@
where the constants $\lambda_{iou}=2$ and $\lambda_{L_{1}}=5$ are the same as in [9, 10], $\gamma$ is set to 5. Our framework is trained end-to-end with the overall loss $\mathcal{L}_{overall}$, after loading the pre-trained weights of the ViT for image classification.
@@ 4 Experiments
In this section, we extensively evaluate our method on five tracking benchmarks, including comparison with state-of-the-art methods, comprehensive ablation studies, and an analysis of generality.
@@ 4.1 Implementation Details
We employ A-ViT-T [39], DeiT-Tiny[33], and ViT-Tiny[33] as the backbones to build our trackers: ABTrack-AViT, ABTrack-DeiT, and ABTrack-ViT. We use the training splits from several datasets for our training process, including GOT-10k [44], LaSOT [45], COCO [46], and TrackingNet [47]. In our experiments, we set the bypass probability threshold $\rho$, block sparsity parameter $\tau_{0}$, $\zeta$, pruning ratio $\mu$, and regularization sparsity hyper-parameter $\alpha_{k}$ to 0.5, 0.4, 0.1, 0.3, and $10^{-4}$, respectively. These parameters were carefully chosen to balance the trade-offs between model complexity, computational efficiency, and tracking performance. We use the AdamW optimizer [48] with a weight decay of $10^{-4}$ for finally training the model, the initial learning rate is set to $4\times 10^{-5}$, and the models are trained for 300 epochs with a batch size of 32. The learning rate is reduced by a factor of 10 after 240 epochs.
@@ 4.2 Datasets
GOT-10k [44] is a large-scale benchmark for generic object tracking in the wild. It contains over 10,000 video sequences with diverse challenges. TrackingNet [47] is a high-quality benchmark for object tracking. It consists of 30,000 video sequences with precise pixel-level annotations, covering a wide range of object categories and tracking scenarios. LaSOT [45] is a comprehensive benchmark for single object tracking. It contains 1,400 video sequences with a total of 3.5 million frames, covering various challenges. UAV123 [49] is a large-scale benchmark for aerial tracking, comprising 123 sequences with over 112K frames. It covers a variety of challenging aerial scenarios. VOT-ST2021 [50] is a visual tracking challenge that tests algorithms on short-term, single object tracking in RGB images.